# The Muckenhoupt condition

### v2

@@ 1.1. The Muckenhoupt condition beyond weighted Lebesgue spaces
$[w]_{p}:=\sup_{Q}\Big{(}\frac{1}{|Q|}\int_{Q}\!w^{p}\,\mathrm{d}x\Big{)}^{\frac{1}{p}}\Big{(}\frac{1}{|Q|}\int_{Q}\!w^{-p^{\prime}}\,\mathrm{d}x\Big{)}^{\frac{1}{p^{\prime}}}<\infty,$
@@
In this case we write $X\in A$ and denote the smallest possible constant $C$ by $[X]_{A}$. Note that for $1\leq p\leq\infty$ and $w$ a weight we have $L^{p}_{w}(\mathbf{R}^{d})\in A$ if and only if $w\in A_{p}$, with
$[L^{p}_{w}(\mathbf{R}^{d})]_{A}=[w]_{p}.$
@@
[…]
@@ 1.2. Generalizing the Muckenhoupt condition
We let $[X]_{A_{\text{strong}}}$ denote the smallest possible $C$ in the above inequality. This condition reduces back to the Muckenhoupt condition $X\in A$ if one only considers collections $\mathcal{P}$ consisting of single cubes. Moreover, as the operator inside the norm on the left-hand side of (1.3) is dominated by $Mf$, this condition is weaker than $M:X\to X$. Nevertheless, Diening shows that if $X=L^{p(\cdot)}(\mathbf{R}^{d})$ and (1.2) holds, then $X\in A_{\text{strong}}$ is equivalent to $M:X\to X$.
@@
The smallest possible constant is denoted by $[X]_{\mathcal{G}}$. This property holds for (weighted) Lebesgue spaces by Hölder’s inequality. Further examples include (weighted) variable Lebesgue spaces with exponent functions satisfying global $\log$-Hölder continuity. We refer the reader to Section 5 for an overview.
@@
$[X]_{A}\leq\|M\|_{X\to X_{\text{weak}}}\lesssim_{d}[X]_{A_{\text{strong}}}\leq\|M\|_{X\to X}.$
@@
$[X]_{A_{\text{strong}}}\leq[X]_{\mathcal{G}}[X]_{A}.$
This result is proven in Theorem 3.14 and Corollary 3.16 below. We also give a new characterization of the property $X\in\mathcal{G}$, generalizing the ideas of [DHHR11, Section 7.3].
@@
Property ii can be interpreted as being able to reconstitute the norm of $f$ through only its local norm behavior. The proof of this result is inspired by the work of Kopaliani in [Kop04], and can be found as part of Theorem 3.9 below.
@@ 1.3. Duality of the Hardy-Littlewood maximal operator
(Q) Given a class of spaces X, when is it true that if M:XX, thenalso $\begin{array}[]{c}\text{\emph{Given a class of spaces $X$, when is it true that if $M:X\to X$, then}}\\ \text{\emph{also $M:X^{\prime}\to X^{\prime}$?}}\end{array}$
@@
$[v]_{\text{FW}}:=\sup_{Q}\frac{1}{v(Q)}\int_{Q}\!M(v\operatorname{\mathbf{1}}_{Q})\,\mathrm{d}x<\infty.$
@@
In general, to answer (Q), we need to impose some kind of reflexivity condition on $X$. Following the above examples, a possible candidate are the spaces $X$ that are $r$-convex and $s$-concave for some $1<r\leq s<\infty$. We note here that all quasi-Banach function spaces are $\infty$-concave by the ideal property, and a quasi-Banach function space is $1$-convex if and only if it is isomorphic to a Banach function space, i.e., there is an equivalent quasinorm on the space that satisfies the triangle inequality. A version of the definition of $r$-convexity and $s$-concavity in the setting of Banach lattices dates back to Krivine and Maurey [Kri74, Mau74], see [CT86, Kal84] for the case of quasi–Banach lattices, and many of the properties of these spaces are explored in these works. See also Subsection 2.3 below.
@@
A self-improvement condition like the one for the bound $M:X\to X$ does not exist for the condition $X\in A_{\text{strong}}$. This is exemplified by the fact that $L^{1}(\mathbf{R}^{d})\in A_{\text{strong}}$, but $L^{1}(\mathbf{R}^{d})^{r}=L^{\frac{1}{r}}(\mathbf{R}^{d})\notin A_{\text{strong}}$ for any $r>1$. However, a self-improvement result was established by Diening for Musielak-Orlicz spaces in [Die05, Theorem 5.7]. Given a Young function $\phi$ satisfying
@@ 2.2. Quasi-Banach function spaces
equipped with $\|\cdot\|_{X^{\prime}}$ is called the Köthe dual of $X$. If $X$ is a Banach function space over $\Omega$, then so is $X^{\prime}$. However, $X^{\prime}$ might not satisfy the saturation property if $X$ is a quasi-Banach function space over $\Omega$, e.g., $X=L^{p}(\mathbf{R}^{d})$ with $10<p<$ 1 satisfies $X^{\prime}=\{0\}$.
@@
The Köthe dual of a quasi-Banach function space satisfies the Fatou property by the monotone convergence theorem. By the Lorentz-Luxemburg theorem, a Banach function space $X$ over $\Omega$ satisfies the Fatou property if and only if $X^{\prime\prime}=X$. Moreover, order-continuity is equivalent to the canonical embedding $X^{\prime}\hookrightarrow X^{\ast}$ being an isomorphism. In particular, $X$ is reflexive if and only if $X$ satisfies the Fatou property and $X$ and $X^{\prime}$ are order-continuous, see [LN24a, Corollary 3.16]. Both the Fatou property and order-continuity are sufficient conditions for the equality
@@ 2.4. Factorization
For any $10<\theta<$ 1 and pair of quasi-Banach function spaces $X_{0}$, $X_{1}$ over $\Omega$, we call
@@
(2.1) $(X^{\theta})^{\prime}=(X^{\prime})^{\theta}\cdot L^{1}(\mathbf{R}^{d})^{1-\theta}=(X^{\prime})^{\theta}\cdot L^{\frac{1}{1-\theta}}(\mathbf{R}^{d}).$
@@ 2.5. Mixed-norm spaces
Moreover, he showed in [Buk87] (see also [Mal89]) that if $X_{0},X_{1},Y_{0},Y_{1}$ are Banach function spaces and $Y_{0}$, $Y_{1}$ have the Fatou property, then for all $10<\theta<$ 1
@@ 3.1. The Muckenhoupt condition
This equivalence is already contained in, e.g., [Ber99], and we include a proof here.
@@
$[X]_{A}:=\sup_{Q}|Q|^{-1}\|\operatorname{\mathbf{1}}_{Q}\|_{X}\|\operatorname{\mathbf{1}}_{Q}\|_{X^{\prime}}<\infty.$
@@
for any cube $Q$, we have $[X]_{A}\geq 1$. Moreover, since the Muckenhoupt condition requires that $\operatorname{\mathbf{1}}_{Q}\in X$ and $\operatorname{\mathbf{1}}_{Q}\in X^{\prime}$ for all cubes $Q$, any space $X\in A$ has the property that $X^{\prime}$ is saturated and, hence, is a Banach function space over $\mathbf{R}^{d}$. Indeed, this follows from the fact that we can pick a sequence of cubes that increases to $\mathbf{R}^{d}$ combined with [LN24a, Proposition 2.5(ii)].
@@ 3.2. Generalized Muckenhoupt conditions
Several further operators related to the boundedness of $M$ and the Muckenhoupt condition with respect to more general averaging operators have appeared in the literature. Given a collection of cubes $\mathcal{P}$ and $f\in L^{1}_{\text{loc}}(\mathbf{R}^{d})$, we define
@@
For $10<\eta<$ 1 we say that a collection of cubes $\mathcal{S}$ is $\eta$-sparse if there exists a pairwise disjoint collection $(E_{Q})_{Q\in\mathcal{S}}$ of subsets $E_{Q}\subseteq Q$ for which $|E_{Q}|\geq\eta|Q|$. We will call $\mathcal{S}$ sparse if it is $\tfrac{1}{2}$-sparse.
@@
    The smallest possible constant $C$ is denoted by $[X]_{A_{\text{strong}}}$.
@@
    The smallest possible constant $C$ is denoted by $[X]_{A_{\text{sparse}}}$.
Since for pairwise disjoint collections $\mathcal{P}$ and $10<\theta<$ 1 we have
@@
$[X^{\theta}]_{A}\leq[X]_{A}^{\theta},\quad[X^{\theta}]_{A_{\text{strong}}}\leq[X]^{\theta}_{\text{strong}}.$
By Theorem 4.2 below, a variant of this result also holds for the condition $X\in A_{\text{sparse}}$.
@@
$[X]_{A}\leq[X]_{A_{\text{strong}}}\leq[X]_{A_{\text{sparse}}}.$
In the definition of $A_{\text{sparse}}$, it suffices to check the result for sparse collections contained in a dyadic grid as follows from the $3^{d}$-lattice theorem combined with the fact that the sparsity constant $\tfrac{1}{2}$ could be replaced by any other $10<\eta<$ 1. For the definition of a dyadic grid and related properties we refer the reader to [LN18]. This latter statement about the sparsity constant follows from the following result.
@@
Let $\mathcal{D}$ be a dyadic grid, let $10<\eta<$ 1, and let $\mathcal{S}\subseteq\mathcal{D}$ be a finite $\eta$-sparse collection. Then for all $10<\nu<$ 1 and all $f\in L^{1}_{\text{loc}}(\mathbf{R}^{d})$ there exists a $\nu$-sparse collection $\mathcal{E}\subseteq\mathcal{D}$ such that
@@
The saturation property follows from the fact that we can pick a sequence of cubes that increases to $\mathbf{R}^{d}$ and [LN24a, Proposition 2.5(ii)].
For the next assertion, let $f\in X$, $g\in X^{\prime}$, and $\mathcal{P}$ a collection of cubes. Then
@@
We define $[X]_{\mathcal{G}}$ as the smallest possible constant $C$ in the above inequality.
@@
$[X]_{A_{\text{strong}}}\leq[X]_{\mathcal{G}}[X]_{A}.$
@@
In the setting of $\mathbf{R}_{+}=[0,\infty)$ rather than $\mathbf{R}^{d}$, it was shown in [Kop04] that if (3.5) holds with the collection $\mathcal{P}$ replaced by any collection of pairwise disjoint sets, then we must be in the above situation where $X=L^{r}_{w}(\mathbf{R}_{+})$ for some weight $w$. However, non-trivial examples of when (3.5) holds were also provided. Indeed, under a modified version of (3.5) to the space $[0,1]$, this was shown to hold for the variable Lebesgue space $X=L^{p(\cdot)}([0,1])$ with exponents $p(\cdot)$ satisfying
@@
An explicit counterexample in this setting is also constructed. We discuss further examples in Section 5.2. The proof uses a Banach function space analogue of the average $\langle f\rangle_{1,Q}$. Indeed, defining
@@
$\langle f\rangle_{1,Q}\leq C\|f\|_{X_{Q}},$
if and only if $X\in A$, where the smallest possible $C$ is given by $[X]_{A}$. Indeed, this follows from the observation that the above inequality is just a rewrite of the boundedness of the averaging operator $T_{Q}$, combined with Proposition 3.3.
@@
If $X\in A$, then $\Rightarrow$ $X\in A_{\text{strong}}$, with
$[X]_{A_{\text{strong}}}\leq C_{3}[X]_{A}.$
@@
Thus, $X\in A_{\text{strong}}$ with $[X]_{A_{\text{strong}}}\leq C_{3}[X]_{A}$, as desired. ∎
@@
$[X^{r}]^{\frac{1}{r}}_{A_{\text{sparse}}}\lesssim_{d,K_{X}}[X]_{A_{\text{sparse}}}.$
@@
$[X^{r}]^{\frac{1}{r}}_{A_{\text{sparse}}}\lesssim_{d,K_{X}}[X]_{A_{\text{sparse}}},$

### v3

@@ 1.1. The Muckenhoupt condition beyond weighted Lebesgue spaces
$w]_{p}:=\sup_{Q}\Big{(}\frac{1}{|Q|}\int_{Q}\!w^{p}\,\mathrm{d}x\Big{)}^{\frac{1}{p}}\Big{(}\frac{1}{|Q|}\int_{Q}\!w^{-p^{\prime}}\,\mathrm{d}x\Big{)}^{\frac{1}{p^{\prime}}}<\infty,$
@@
In this case we write $X\in A$ and denote the smallest possible constant $C$ by $X]_{A}$. Note that for $1\leq p\leq\infty$ and $w$ a weight we have $L^{p}_{w}(\mathbf{R}^{d})\in A$ if and only if $w\in A_{p}$, with
$L^{p}_{w}(\mathbf{R}^{d})]_{A}=[w]_{p}.$
@@
Alternatively, one can also prove the implications $i\Rightarrowi&i$v through Rubio de Francia extrapolation. Early works on extrapolation in Banach function spaces date back to [CGMP06]. The argument for our implication can already be found in [CMP11], and was stated explicitly in [Cru17, Theorem 10.1], albeit for function spaces satisfying the definition of [BS88] (see [LN24a] for a thorough discussion on the appropriate assumptions to make in the definition of a Banach function space). Further extrapolation results and comparisons can be found in [CMM22, Nie23].
@@ 1.2. Generalizing the Muckenhoupt condition
We let $X]_{A_{\text{strong}}}$ denote the smallest possible $C$ in the above inequality. This condition reduces back to the Muckenhoupt condition $X\in A$ if one only considers collections $\mathcal{P}$ consisting of single cubes. Moreover, as the operator inside the norm on the left-hand side of (1.3) is dominated by $Mf$, this condition is weaker than $M:X\to X$. Nevertheless, Diening shows that if $X=L^{p(\cdot)}(\mathbf{R}^{d})$ and (1.2) holds, then $X\in A_{\text{strong}}$ is equivalent to $M:X\to X$.
@@
The smallest possible constant is denoted by $X]_{\mathcal{G}}$. This property holds for (weighted) Lebesgue spaces by Hölder’s inequality. Further examples include (weighted) variable Lebesgue spaces with exponent functions satisfying global $\log$-Hölder continuity. We refer the reader to Section 5 for an overview.
@@
$X]_{A}\leq\|M\|_{X\to X_{\text{weak}}}\lesssim_{d}[X]_{A_{\text{strong}}}\leq\|M\|_{X\to X}.$
@@
$X]_{A_{\text{strong}}}\leq[X]_{\mathcal{G}}[X]_{A}.$
This result is proved in Theorem 3.15 and Corollary 3.17 below. We also give a new characterization of the property $X\in\mathcal{G}$, generalizing the ideas of [DHHR11, Section 7.3].
@@
Property $i$ can be interpreted as being able to reconstitute the norm of $f$ through only its local norm behavior. The proof of this result is inspired by the work of Kopaliani in [Kop04], and can be found as part of Theorem 3.9 below.
@@ 1.3. Duality of the Hardy-Littlewood maximal operator
[…]
@@
$v]_{\text{FW}}:=\sup_{Q}\frac{1}{v(Q)}\int_{Q}\!M(v\operatorname{\mathbf{1}}_{Q})\,\mathrm{d}x<\infty.$
@@
In general, to answer (Q), we need to impose some kind of reflexivity condition on $X$. Looking at the above examples, one could expect that the condition that $X$ is $r$-convex and $s$-concave for some $1<r\leq s<\infty$ is a suitable choice. We note here that all quasi-Banach function spaces are $\infty$-concave by the ideal property, and a quasi-Banach function space is $1$-convex if and only if it is isomorphic to a Banach function space, i.e., there is an equivalent quasinorm on the space that satisfies the triangle inequality. A version of the definition of $r$-convexity and $s$-concavity in the setting of Banach lattices dates back to Krivine and Maurey [Kri74, Mau74], see [CT86, Kal84] for the case of quasi–Banach lattices, and many of the properties of these spaces are explored in these works. See also Subsection 2.3 below.
@@
A self-improvement condition like the one for the bound $M:X\to X$ does not exist for the condition $X\in A_{\text{strong}}$. This is exemplified by the fact that $L^{1}(\mathbf{R}^{d})\in A_{\text{strong}}$, but $L^{1}(\mathbf{R}^{d})^{r}=L^{\frac{1}{r}}(\mathbf{R}^{d})\notin A_{\text{strong}}$ for any $r>1$. However, a self-improvement result was established by Diening for Musielak-Orlicz spaces in [Die05, Theorem 5.7]. Given a $\Phi$-function $\phi$ satisfying
@@ 2.2. Quasi-Banach function spaces
equipped with $\|\cdot\|_{X^{\prime}}$ is called the Köthe dual of $X$. If $X$ is a Banach function space over $\Omega$, then so is $X^{\prime}$. However, $X^{\prime}$ might not satisfy the saturation property if $X$ is a quasi-Banach function space over $\Omega$, e.g., $X=L^{p}(\mathbf{R}^{d})$ with $0<p<1$ satisfies $X^{\prime}=\{0\}$.
In the case that $\Omega=\mathbf{R}^{d}$, an often used sufficient condition for the saturation property is as follows.
Proposition 2.2.
Let $X\subseteq L^{0}(\mathbf{R}^{d})$ be a complete quasi-normed vector-space satisfying the ideal property. If $\operatorname{\mathbf{1}}_{Q}\in X$ for all cubes $Q\subseteq\mathbf{R}^{d}$, then $X$ satisfies the saturation property and, hence, is a quasi-Banach function space over $\mathbf{R}^{d}$.
Proof.
By [LN24a, Proposition 2.5(ii)], the saturation property is equivalent to the existence of a sequence of sets increasing to $\mathbf{R}^{d}$ whose indicator functions belong to $X$. Thus, we can pick any sequence of cubes that increases to $\mathbf{R}^{d}$ to satisfy this. ∎
@@
The Köthe dual of a quasi-Banach function space satisfies the Fatou property by the monotone convergence theorem. By the Lorentz-Luxemburg theorem, a Banach function space $X$ over $\Omega$ satisfies the Fatou property if and only if it is reflexive in the sense of Köthe duality. This result can be found, e.g., in [Zaa67, Theorem 71.1].
Theorem (Lorentz-Luxemburg).
Let $X$ be a Banach function space over $\Omega$. Then $X$ satisfies the Fatou property if and only if $X^{\prime\prime}=X$.
Order-continuity is equivalent to the canonical embedding $X^{\prime}\hookrightarrow X^{\ast}$ being an isomorphism. In particular, $X$ is reflexive if and only if $X$ satisfies the Fatou property and $X$ and $X^{\prime}$ are order-continuous, see [LN24a, Corollary 3.16]. Both the Fatou property and order-continuity are sufficient conditions for the equality
@@ 2.4. Factorization
For any $0<\theta<1$ and pair of quasi-Banach function spaces $X_{0}$, $X_{1}$ over $\Omega$, we call
@@
(2.1) $(X^{\theta})^{\prime}=(X^{\prime})^{\theta}\cdot L^{1}(\Omega)^{1-\theta}=(X^{\prime})^{\theta}\cdot L^{\frac{1}{1-\theta}}(\Omega).$
@@ 2.5. Mixed-norm spaces
Moreover, he showed in [Buk87] (see also [Mal89]) that if $X_{0},X_{1},Y_{0},Y_{1}$ are Banach function spaces and $Y_{0}$, $Y_{1}$ have the Fatou property, then for all $0<\theta<1$
@@ 3.1. The Muckenhoupt condition
This equivalence is already contained in, e.g., [Ber99], but we include a proof here.
@@
$X]_{A}:=\sup_{Q}|Q|^{-1}\|\operatorname{\mathbf{1}}_{Q}\|_{X}\|\operatorname{\mathbf{1}}_{Q}\|_{X^{\prime}}<\infty.$
@@
for any cube $Q$, we have $X]_{A}\geq 1$. Moreover, since the Muckenhoupt condition requires that $\operatorname{\mathbf{1}}_{Q}\in X$ and $\operatorname{\mathbf{1}}_{Q}\in X^{\prime}$ for all cubes $Q$, any space $X\in A$ has the property that $X^{\prime}$ is saturated and, hence, is a Banach function space over $\mathbf{R}^{d}$ by Proposition 2.2.
@@ 3.2. Generalized Muckenhoupt conditions
Two other more general averaging operators related to the boundedness of $M$ and the Muckenhoupt condition have appeared in the literature. Given a collection of cubes $\mathcal{P}$ and $f\in L^{1}_{\text{loc}}(\mathbf{R}^{d})$, we define
@@
For $0<\eta<1$ we say that a collection of cubes $\mathcal{S}$ is $\eta$-sparse if there exists a pairwise disjoint collection $(E_{Q})_{Q\in\mathcal{S}}$ of subsets $E_{Q}\subseteq Q$ for which $|E_{Q}|\geq\eta|Q|$. We will call $\mathcal{S}$ sparse if it is $\tfrac{1}{2}$-sparse.
@@
    The smallest possible constant $C$ is denoted by $X]_{A_{\text{strong}}}$.
@@
    The smallest possible constant $C$ is denoted by $X]_{A_{\text{sparse}}}$.
Since for pairwise disjoint collections $\mathcal{P}$ and $0<\theta<1$ we have
@@
$X^{\theta}]_{A}\leq[X]_{A}^{\theta},\quad[X^{\theta}]_{A_{\text{strong}}}\leq[X]^{\theta}_{\text{strong}}.$
By combining Theorem 4.2 and Theorem 3.15 below, a variant of this result also holds for the condition $X\in A_{\text{sparse}}$. More precisely, if $X\in A_{\text{sparse}}$ and $0<\theta<1$, then $X^{\theta}\in A_{\text{sparse}}$ with
$X^{\theta}]_{A_{\text{sparse}}}\lesssim_{d}\frac{1}{\theta}\frac{1}{1-\theta}[X]_{A_{\text{sparse}}}^{1+\theta}.$
@@
$X]_{A}\leq[X]_{A_{\text{strong}}}\leq[X]_{A_{\text{sparse}}}.$
In the definition of $A_{\text{sparse}}$, it suffices to check the result for sparse collections contained in a dyadic grid as follows from the $3^{d}$-lattice theorem combined with the fact that the sparsity constant $\tfrac{1}{2}$ could be replaced by any other $0<\eta<1$. For the definition of a dyadic grid and related properties we refer the reader to [LN18]. This latter statement about the sparsity constant follows from the following result.
@@
Let $\mathcal{D}$ be a dyadic grid, let $0<\eta<1$, and let $\mathcal{S}\subseteq\mathcal{D}$ be a finite $\eta$-sparse collection. Then for all $0<\nu<1$ and all $f\in L^{1}_{\text{loc}}(\mathbf{R}^{d})$ there exists a $\nu$-sparse collection $\mathcal{E}\subseteq\mathcal{D}$ such that
@@
The saturation property follows from Proposition 2.2. For the next assertion, let $f\in X$, $g\in X^{\prime}$, and $\mathcal{P}$ a collection of cubes. Then
@@
We define $X]_{\mathcal{G}}$ as the smallest possible constant $C$ in the above inequality.
@@
$X]_{A_{\text{strong}}}\leq[X]_{\mathcal{G}}[X]_{A}.$
@@
In the setting of $\mathbf{R}_{+}=[0,\infty)$ rather than $\mathbf{R}^{d}$, it was shown in [Kop04] that if (3.5) holds with the collection $\mathcal{P}$ replaced by any collection of pairwise disjoint sets, then we must be in the above situation where $X=L^{r}_{w}(\mathbf{R}_{+})$ for some weight $w$. However, non-trivial examples of when (3.5) holds were also provided. Indeed, under the condition (3.5) specialized to the space $[0,1]$ rather than $\mathbf{R}^{d}$, this was shown to hold for the variable Lebesgue space $X=L^{p(\cdot)}([0,1])$ with exponents $p(\cdot)$ satisfying
@@
An explicit example of a space in this setting that does not satisfy (3.5) is also constructed. We discuss further examples in Section 5.2. The proof uses a Banach function space analogue of the average $\langle f\rangle_{1,Q}$. Indeed, defining
@@
$\langle f\rangle_{1,Q}\leq C\|f\|_{X_{Q}}$
if and only if $X\in A$, where the smallest possible $C$ is given by $X]_{A}$. Indeed, this follows from the observation that the above inequality is just a rewrite of the boundedness of the averaging operator $T_{Q}$, combined with Proposition 3.3.
@@
[…]
$X]_{A_{\text{strong}}}\leq C_{3}[X]_{A}.$
@@
Thus, $X\in A_{\text{strong}}$ with $X]_{A_{\text{strong}}}\leq C_{3}[X]_{A}$, as desired. ∎
@@
Remark 3.11.
Theorem 3.9c gives a sufficient condition under which $A$ and $A_{\text{strong}}$ are equivalent. This leads one to wonder if there is such a condition that characterizes when $A$ and $A_{\text{strong}}$ are equivalent. More precisely, one can ask whether there is a suitably nice condition $\mathcal{H}$ for which $X\in A\cap\mathcal{H}$ is equivalent to $X\in A_{\text{strong}}$ for all Banach function spaces $X$ over $\mathbf{R}^{d}$ with the Fatou property. We leave this problem as an open question.
@@
$X^{r}]^{\frac{1}{r}}_{A_{\text{sparse}}}\lesssim_{d,K_{X}}[X]_{A_{\text{sparse}}}.$
@@
$X^{r}]^{\frac{1}{r}}_{A_{\text{sparse}}}\lesssim_{d,K_{X}}[X]_{A_{\text{sparse}}},$